# FAST: Foreground‑aware Diffusion with Accelerated Sampling Trajectory for Segmentation‑oriented Anomaly Synthesis

### v4

@@ 1 Introduction
Limitations. Despite recent advances, current anomaly synthesis methods face three fundamental limitations that hinder their effectiveness for segmentation tasks xu2025survey. (i) Lack of controllability. Most existing methods provide limited control over the structure, location, or extent of synthesized anomalies. This limitation is particularly evident in GAN-based approaches niu2020defect; zhang2021defect; du2022new. These methods typically adopt a one-shot generation paradigm, offering little flexibility in specifying where and how anomalies should appear. (ii) Neglect of segmentation-relevant properties. Training-free methods such as patch replacement or texture corruption li2021cutpaste; zavrtanik2021draem may produce visible anomalies, but the synthesized patterns often lack the structural consistency and complexity of real-world industrial anomalies, which are critical for improving segmentation performance. (iii) Uniform treatment of spatial regions and inefficiency. Although recent diffusion-based methods he2024anomalycontrol; hu2024anomalyxfusion; shi2025few have mitigated the above issues, they still treat all spatial regions uniformly during both forward and reverse processes, without explicitly modeling the distinct statistical properties of anomaly regions zhang2024realnet; yao2024glad. This absence of region-aware modeling prevents the model from preserving abnormal regions throughout the synthesis trajectory. Moreover, these models typically require hundreds to thousands of denoising steps ho2020denoising; song2021denoising, resulting in a significant computational cost, especially for the real-world production line changeover. While recent training-free methods liu2022pseudo aim to accelerate sampling, they fail to incorporate anomaly-aware cues, making them less effective for segmentation-oriented industrial anomaly synthesis (SIAS). These limitations motivate the need for SIAS models that support controllable anomaly synthesis, explicit modeling of anomaly regions, and efficient, task-aligned sampling strategies.
@@ 2 Related work
Industrial Anomaly Synthesis. Industrial anomaly synthesis aims to mitigate the scarcity of labeled abnormal samples in real-world inspection scenarios. Existing methods can be categorized into hand-crafted and DL-based approaches. Hand-crafted methods typically apply training-free manipulations to normal images, such as patch pasting pei2023self; schluter2022natural or external texture blending zavrtanik2021draem; yang2023memseg; zhang2023destseg from sources like DTD cimpoi14describing, but they suffer from distributional deviation and limited realism. Deep learning-based methods alleviate these limitations by learning from real anomaly patterns. GAN-based methods duan2023few; wen2022new can synthesize visually realistic anomalies but lack fine-grained controllability over anomaly shape and location. Diffusion-based methods duan2023few; hu2024anomalydiffusion; yang2024defect; gui2024few offer stronger generative capacity via large-scale pretrained models, yet treat all regions uniformly and lack explicit control over anomaly localization, which is essential for segmentation. To this end, we propose FAST, which integrates foreground-aware reconstruction and efficient, segmentation-oriented anomaly synthesis into a unified diffusion framework.
@@
Acceleration of Discrete-Time Diffusion Models. Diffusion models can be categorized into continuous-time and discrete-time frameworks. Continuous formulations lu2022dpm; lu2022dpm2; zheng2025diffusion adopt SDE/ODE-based parameterizations and leverage high-order solvers for efficient sampling. In contrast, standard DDPMs ho2020denoising model a discrete-time Markov chain with fixed variance schedules and require thousands of iterative denoising steps. While continuous-time solvers achieve notable speedups, they rely on continuously parameterized noise or score functions, which requires reformulating training objectives or interface in discrete-time models. Therefore, various acceleration techniques have been developed specifically for discrete-time diffusion. Some methods modify the generative process to reduce steps: DDGAN xiao2022tackling integrates GAN-based decoding, TLDM zheng2023truncated and ES-DDPM lyu2022accelerating truncate the forward process, and Blurring Diffusion Models hoogeboom2023blurring operate in the frequency domain. However, these methods require retraining and show limited generalization. In contrast, training-free approaches such as DDIM song2021denoising, PLMS liu2022pseudo, and GGDM watson2022learning accelerate sampling without model modification. Yet, they treat all spatial regions uniformly and lack task-specific guidance essential for SIAS. Recent work like CUT sun2024cut introduces external prompts for localized control for anomalies, but at the cost of multiple iterations per sampling step. In comparison, FAST proposes a novel training-free strategy that aggregates multiple denoising steps into coarse-to-fine segments while injecting mask-aware structural guidance, enabling efficient SIAS.
@@
Foreground–background Decoupling. Foreground–background decoupling has been widely employed in industrial anomaly synthesis to enhance spatial precision and suppress irrelevant background interference. The core idea is to isolate defect-related regions from normal contexts, thereby improving inspection accuracy and synthesis controllability. Most methods such as PRN prn and DCDGANc wei2023diversified perform explicit two-stage compositions, which first generate abnormal foregrounds and then blend them with normal backgrounds under soft mask constraints, but often suffer from boundary inconsistencies. Recent studies have introduced implicit separation; for instance, FCIS wang2025enhanced enlarges the anomaly–background distance via contrastive learning, while BDG cnmsbsqj incorporates masked attention and regularization within the denoiser to disentangle the influence of anomalies from the surrounding background. Although both BDG and FAST involve diffusion-based synthesis with certain forms of foreground–background decoupling, they pursue different research objectives through fundamentally distinct methodologies. FAST is a segmentation-oriented anomaly synthesis framework that emphasizes pixel-wise structural alignment and contextual consistency, whereas BDG primarily targets robust anomaly detection. Technically, AIAS in FAST analytically aggregates multiple DDPM reverse transitions into a few closed-form, coarse-to-fine updates, forming a deterministic and training-free sampler (e.g., $x_{t}\rightarrow x_{t-1}$) whose coefficients are precomputed under the original variance schedule, without any variance-controlling parameters like DDIM song2021denoising. In contrast, BDG depends on DDIM inversion (e.g., $x_{t-1}\rightarrow x_{t}$) to maintain background features, which requires inversion consistency and retraining with regularization losses. These two mechanisms are fundamentally distinct and not directly interchangeable. Furthermore, the FARM module in FAST functions as an external foreground-reconstruction pathway that injects anomaly-aware noise via masks across timesteps to preserve anomaly salience throughout the sampling trajectory, whereas BDG employs masks merely as internal attention gates to localize edits within the denoiser. Essentially, BDG modifies the attention dynamics inside the denoiser to limit interference, while FARM operates outside the denoiser as a reconstruction-based enhancement module. These conceptual and algorithmic distinctions, together with different experiments and evaluation (segmentation-oriented mIoU/Acc vs. detection-oriented AUROC/AP) demonstrate that FAST and BDG follow independent research lines and remain technically and theoretically original.
@@ 3 Methods
FAST for Anomaly Segmentation. The proposed FAST framework is built upon the LDM rombach2022high of $T$ steps. For notational simplicity, we denote the encoded latent of the original image as $x_{0}$, and its predicted reconstruction from the network as $\hat{x}_{0}$. We define $x_{t_{s}}$ as the noisy latent at timestep $t_{s}$, and $\hat{x}_{t_{s}}$ as the FARM-adjusted, anomaly-aware latent at the same step. Let $\mathcal{M}\in\{0,1\}^{H\times W}$ denote the binary anomaly mask, and $t_{s},t_{e}]$ represent a coarse-to-fine segment in AIAS, where $t_{e}<t_{s}$. Fig. 1 illustrates a single forward-reverse process at step $t_{s}$. In the forward phase, noise is added up to timestep $t_{s}$, yielding a noisy latent $x_{t_{s}}$. FARM ($F_{\phi}$ in Algorithm 1) then predicts a pseudo-clean anomaly latent $\hat{x}_{0}^{\text{an}}$, and adds noise to it up to timestep $t_{s}$ to obtain an anomaly-aware latent $\hat{x}_{t_{s}}$, which aims to match the observed $x_{t_{s}}$ in masked regions during training. In the corresponding reverse process, we divide the full denoising process into $S$ segments, each spanning $t_{s},t_{e}]$. Within each segment, AIAS approximates the posterior transition using: $q(x_{t_{e}}\mid x_{t_{s}},\hat{x}_{0})$. This formulation aggregates multiple DDPM steps into a single numerical update. FARM is also applied to refine $x_{t_{e}}$, ensuring the preservation of anomaly cues throughout the reverse process. More details can be seen in Algorithms 1 and 2.In addition, for the textual conditioning component of LDM, we follow the configuration of Anomaly Diffusion hu2024anomalydiffusion; more implementation details can be found there.
@@ 4 Experiments
Datasets. We evaluate FAST on two widely-used industrial anomaly segmentation benchmarks: MVTec-AD bergmann2019mvtec and BTAD btad. For each anomaly class, we synthesize image–mask pairs using normal images, binary masks, and text prompts describing anomaly semantics. A total of 500 samples are generated for each anomaly type within a class, with approximately one-third used for training and the remainder reserved for evaluation. This design ensures sufficient structural diversity while maintaining training efficiency. Mask Generation Strategy. Our mask synthesis consists of two complementary components: (i) geometric augmentation of real anomaly masks via operations like rotation and flipping; (ii) synthesis of new masks using a Latent Diffusion Model (LDM) pre-trained on real anomaly mask examples, which follows the protocol of AnomalyDiffusion hu2024anomalydiffusion. All synthesized masks undergo manual screening to guarantee visual realism, structural diversity, and consistency with typical industrial abnormal structures. Evaluation Metrics. We report performance using mean intersection over union (mIoU) and pixel-wise accuracy (Acc), following standard practice in anomaly segmentation. Baselines. FAST is compared against six representative anomaly synthesis approaches: CutPaste li2021cutpaste, DRAEM zavrtanik2021draem, GLASS glass, the GAN-based SOTA method DFMGAN dfmgan, and diffusion-based SOTA models Anomaly Diffusion hu2024anomalydiffusion and RealNet zhang2024realnet. To simulate realistic deployment scenarios, we pair all generation methods with lightweight segmentation networks, including Segformer xie2021segformer, BiSeNet V2 yu2021bisenet, and STDC stdc. As our method adopts the same prompt-driven synthesis setup as AnomalyDiffusion hu2024anomalydiffusion, we omit the details here for brevity. Full specifications of the textual configuration, as well as other implementation details, including dataset preprocessing, sampling schedules, loss weights, and hyperparameter settings, are provided in the Supplementary Materials A.5.
@@ 4.3 Ablation Studies
The Impact of AIAS. We compare our proposed AIAS strategy with several widely-used training-free samplers, including DDPM ho2020denoising with 1000 steps, DDIM song2021denoising with 50 steps and PLMS liu2022pseudo with 50 steps. These methods represent state-of-the-art discrete-time sampling approaches for diffusion-based models. To ensure fairness, we exclude continuous-time solvers, as they rely on a fundamentally different formulation based on ODEs or SDEs, which necessitates a distinct training paradigm and architectural adjustments incompatible with our discrete-time framework. Quantitative results are reported in Fig. 5. While DDPM achieves competitive results on certain categories (e.g., carpet, tile), it requires 1000 iterative steps, making it over 20× slower than AIAS in practice.
@@ A.1 Proof of Lemma 1
  • Assume Eq. 14 and .15 hold for $m=r$ with $1\leq r<k$:
@@ A.5 Training Configuration
  • Prompt Construction. For the MVTec dataset, prompts are formed by appending the anomaly type to the product category name. For BTAD, due to anonymized category labels, we use a generic prompt: “damaged". Textual embeddings follow the protocol of AnomalyDiffusion, where each prompt is tokenized into 8 discrete units and embedded using a pre-trained BERT encoder [devlin-etal-2019-bert].

### v5

@@ 1 Introduction
Limitations. Despite recent advances, current anomaly synthesis methods face three fundamental limitations that hinder their effectiveness for segmentation tasks Xu et al. (2025). (i) Lack of controllability. Most existing methods provide limited control over the structure, location, or extent of synthesized anomalies. This limitation is particularly evident in GAN-based approaches Niu et al. (2020); Zhang et al. (2021); Du et al. (2022). These methods typically adopt a one-shot generation paradigm, offering little flexibility in specifying where and how anomalies should appear. (ii) Neglect of segmentation-relevant properties. Training-free methods such as patch replacement or texture corruption Li et al. (2021); Zavrtanik et al. (2021) may produce visible anomalies, but the synthesized patterns often lack the structural consistency and complexity of real-world industrial anomalies, which are critical for improving segmentation performance. (iii) Uniform treatment of spatial regions and inefficiency. Although recent diffusion-based methods He et al. (2024); Hu et al. (2024a); Shi et al. (2025) have mitigated the above issues, they still treat all spatial regions uniformly during both forward and reverse processes, without explicitly modeling the distinct statistical properties of anomaly regions Zhang et al. (2024); Yao et al. (2024). This absence of region-aware modeling prevents the model from preserving abnormal regions throughout the synthesis trajectory. Moreover, these models typically require hundreds to thousands of denoising steps Ho et al. (2020); Song et al. (2021), resulting in a significant computational cost, especially for the real-world production line changeover. While recent training-free methods Liu et al. (2022) aim to accelerate sampling, they fail to incorporate anomaly-aware cues, making them less effective for segmentation-oriented industrial anomaly synthesis (SIAS). These limitations motivate the need for SIAS models that support controllable anomaly synthesis, explicit modeling of anomaly regions, and efficient, task-aligned sampling strategies.
@@ 2 Related work
Industrial Anomaly Synthesis. Industrial anomaly synthesis aims to mitigate the scarcity of labeled abnormal samples in real-world inspection scenarios. Existing methods can be categorized into hand-crafted and DL-based approaches. Hand-crafted methods typically apply training-free manipulations to normal images, such as patch pasting Pei et al. (2023); Schlüter et al. (2022) or external texture blending Zavrtanik et al. (2021); Yang et al. (2023); Zhang et al. (2023b) from sources like DTD Cimpoi et al. (2014), but they suffer from distributional deviation and limited realism. Deep learning-based methods alleviate these limitations by learning from real anomaly patterns. GAN-based methods Duan et al. (2023b); Wen et al. (2022) can synthesize visually realistic anomalies but lack fine-grained controllability over anomaly shape and location. Diffusion-based methods Duan et al. (2023b); Hu et al. (2024b); Yang et al. (2024); Gui et al. (2024) offer stronger generative capacity via large-scale pretrained models, yet treat all regions uniformly and lack explicit control over anomaly localization, which is essential for segmentation. To this end, we propose FAST, which integrates foreground-aware reconstruction and efficient, segmentation-oriented anomaly synthesis into a unified diffusion framework.
@@
Acceleration of Discrete-Time Diffusion Models. Diffusion models can be categorized into continuous-time and discrete-time frameworks. Continuous formulations Lu et al. (2022a, b); Zheng et al. (2025) adopt SDE/ODE-based parameterizations and leverage high-order solvers for efficient sampling. In contrast, standard DDPMs Ho et al. (2020) model a discrete-time Markov chain with fixed variance schedules and require thousands of iterative denoising steps. While continuous-time solvers achieve notable speedups, they rely on continuously parameterized noise or score functions, which requires reformulating training objectives or interface in discrete-time models. Therefore, various acceleration techniques have been developed specifically for discrete-time diffusion. Some methods modify the generative process to reduce steps: DDGAN Xiao et al. (2022) integrates GAN-based decoding, TLDM Zheng et al. (2023) and ES-DDPM Lyu et al. (2022) truncate the forward process, and Blurring Diffusion Models Hoogeboom and Salimans (2023) operate in the frequency domain. However, these methods require retraining and show limited generalization. In contrast, training-free approaches such as DDIM Song et al. (2021), PLMS Liu et al. (2022), and GGDM Watson et al. (2022) accelerate sampling without model modification. Yet, they treat all spatial regions uniformly and lack task-specific guidance essential for SIAS. Recent work like CUT Sun et al. (2024) introduces external prompts for localized control for anomalies, but at the cost of multiple iterations per sampling step. In comparison, FAST proposes a novel training-free strategy that aggregates multiple denoising steps into coarse-to-fine segments while injecting mask-aware structural guidance, enabling efficient SIAS.
@@
Foreground–background Decoupling. Foreground–background decoupling has been widely employed in industrial anomaly synthesis to enhance spatial precision and suppress irrelevant background interference. The core idea is to isolate defect-related regions from normal contexts, thereby improving downstream performance and synthesis controllability. Most methods such as PRN Zhang et al. (2023a) and DCDGANc Wei et al. (2023) perform explicit two-stage compositions, which first generate abnormal foregrounds and then blend them with normal backgrounds under soft mask constraints, but often suffer from boundary inconsistencies. Recent studies have introduced implicit separation; for instance, FCIS Wang et al. (2025) enlarges the anomaly–background distance via contrastive learning, while BDG Cho et al. (2024) incorporates masked attention and regularization within the denoiser to disentangle the influence of anomalies from the surrounding background. Although both BDG and FAST involve diffusion-based synthesis with certain forms of foreground–background decoupling, they pursue different research objectives through fundamentally distinct methodologies. FAST is a segmentation-oriented anomaly synthesis framework that emphasizes pixel-wise structural alignment and contextual consistency, whereas BDG primarily targets robust anomaly detection. Technically, AIAS in FAST analytically aggregates multiple DDPM reverse transitions into a few closed-form, coarse-to-fine updates, forming a deterministic and training-free sampler (e.g., $x_{t}\rightarrow x_{t-1}$) whose coefficients are precomputed under the original variance schedule, without any variance-controlling parameters like DDIM Song et al. (2021). In contrast, BDG depends on DDIM inversion (e.g., $x_{t-1}\rightarrow x_{t}$) to maintain background features, which requires inversion consistency and retraining with regularization losses. These two mechanisms are fundamentally distinct and not directly interchangeable. Furthermore, the FARM module in FAST functions as an external foreground-reconstruction pathway that injects anomaly-aware noise via masks across timesteps to preserve anomaly salience throughout the sampling trajectory, whereas BDG employs masks merely as internal attention gates to localize edits within the denoiser. Essentially, BDG modifies the attention dynamics inside the denoiser to limit interference, while FARM operates outside the denoiser as a reconstruction-based enhancement module. These conceptual and algorithmic distinctions, together with different experiments and evaluation (segmentation-oriented mIoU/Acc vs. detection-oriented AUROC/AP) demonstrate that FAST and BDG follow independent research lines and remain technically and theoretically original.
@@ 3 Methods
FAST for Anomaly Segmentation. The proposed FAST framework is built upon the LDM Rombach et al. (2022) of $T$ steps. For notational simplicity, we denote the encoded latent of the original image as $x_{0}$, and its predicted reconstruction from the network as $\hat{x}_{0}$. We define $x_{t_{s}}$ as the noisy latent at timestep $t_{s}$, and $\hat{x}_{t_{s}}$ as the FARM-adjusted, anomaly-aware latent at the same step. Let $\mathcal{M}\in\{0,1\}^{H\times W}$ denote the binary anomaly mask, and $t_{s},t_{e}]$ represent a coarse-to-fine segment in AIAS, where $t_{e}<t_{s}$. Fig. 1 illustrates a single forward-reverse process at step $t_{s}$. In the forward phase, noise is added up to timestep $t_{s}$, yielding a noisy latent $x_{t_{s}}$. FARM ($F_{\phi}$ in Algorithm 1) then predicts a pseudo-clean anomaly latent $\hat{x}_{0}^{\text{an}}$, and adds noise to it up to timestep $t_{s}$ to obtain an anomaly-aware latent $\hat{x}_{t_{s}}$, which aims to match the observed $x_{t_{s}}$ in masked regions during training. In the corresponding reverse process, we divide the full denoising process into $S$ segments, each spanning $t_{s},t_{e}]$. Within each segment, AIAS approximates the posterior transition using: $q(x_{t_{e}}\mid x_{t_{s}},\hat{x}_{0})$. This formulation aggregates multiple DDPM steps into a single numerical update. FARM is also applied to refine $x_{t_{e}}$, ensuring the preservation of anomaly cues throughout the reverse process. More details can be seen in Algorithms 1 and 2.In addition, for the textual conditioning component of LDM, we follow the configuration of Anomaly Diffusion Hu et al. (2024b); more implementation details can be found there.
@@ 3.1 Anomaly-Informed Accelerated Sampling
$\varepsilon^{(m)}_{k}\sim\mathcal{N}\!\bigl(0,\Sigma^{(m)}_{k}\bigr),\qquad\Sigma^{(m)}_{k}=\sum_{i=0}^{m-1}\Bigl(\prod_{j=1}^{i}C_{k-j}\Bigr)\Sigma_{k-i}\Bigl(\prod_{j=1}^{i}C_{k-j}\Bigr)^{\!\top}.$
@@ 4 Experiments
Datasets. We evaluate FAST on two widely-used industrial anomaly segmentation benchmarks: MVTec-AD Bergmann et al. (2019) and BTAD Mishra et al. (2021). For each anomaly class, we synthesize image–mask pairs using normal images, binary masks, and text prompts describing anomaly semantics. A total of 500 samples are generated for each anomaly type within a class, with approximately one-third used for training and the remainder reserved for evaluation. This design ensures sufficient structural diversity while maintaining training efficiency. Mask Generation Strategy. Our mask synthesis consists of two complementary components: (i) geometric augmentation of real anomaly masks via operations like rotation and flipping; (ii) synthesis of new masks using a Latent Diffusion Model (LDM) pre-trained on real anomaly mask examples, which follows the protocol of AnomalyDiffusion Hu et al. (2024b). All synthesized masks undergo manual screening to guarantee visual realism, structural diversity, and consistency with typical industrial abnormal structures. Evaluation Metrics. We report performance using mean intersection over union (mIoU) and pixel-wise accuracy (Acc), following standard practice in anomaly segmentation. Baselines. FAST is compared against six representative anomaly synthesis approaches: CutPaste Li et al. (2021), DRAEM Zavrtanik et al. (2021), GLASS Chen et al. (2025), the GAN-based SOTA method DFMGAN Duan et al. (2023a), and diffusion-based SOTA models Anomaly Diffusion Hu et al. (2024b) and RealNet Zhang et al. (2024). To simulate realistic deployment scenarios, we pair all generation methods with lightweight segmentation networks, including Segformer Xie et al. (2021), BiSeNet V2 Yu et al. (2021), and STDC Fan et al. (2021). As our method adopts the same prompt-driven synthesis setup as AnomalyDiffusion Hu et al. (2024b), we omit the details here for brevity. Full specifications of the textual configuration, as well as other implementation details, including dataset preprocessing, sampling schedules, loss weights, and hyperparameter settings, are provided in the Supplementary Materials A.5.
@@ 4.3 Ablation Studies
The Impact of AIAS. We compare our proposed AIAS strategy with several widely-used training-free samplers, including DDPM Ho et al. (2020) with 1000 steps, DDIM Song et al. (2021) with 50 steps and PLMS Liu et al. (2022) with 50 steps. These methods represent state-of-the-art discrete-time sampling approaches for diffusion-based models. To ensure fairness, we exclude continuous-time solvers, as they rely on a fundamentally different formulation based on ODEs or SDEs, which necessitates a distinct training paradigm and architectural adjustments incompatible with our discrete-time framework. Quantitative results are reported in Fig. 5. While DDPM achieves competitive results on certain categories (e.g., carpet, tile), it requires 1000 iterative steps, making it over 20× slower than AIAS in practice.
@@ A.1 Proof of Lemma 1
  • Assume Eq. 13 and .14 hold for $m=r$ with $1\leq r<k$:
@@ A.5 Training Configuration
  • Prompt Construction. For the MVTec dataset, prompts are formed by appending the anomaly type to the product category name. For BTAD, due to anonymized category labels, we use a generic prompt: “damaged". Textual embeddings follow the protocol of AnomalyDiffusion, where each prompt is tokenized into 8 discrete units and embedded using a pre-trained BERT encoder [5].